# Rank1: Test-Time Compute for Reranking in Information Retrieval

### v1

@@ Abstract
We introduce Rank1, the first reranking model trained to take advantage of test-time compute. Rank1 demonstrates the applicability within retrieval of using a reasoning language model (i.e. OpenAI’s o1, Deepseek’s R1, etc.) for distillation in order to rapidly improve the performance of a smaller model. We gather and open-source a dataset of more than 600,000 examples of R1 reasoning traces from queries and passages in MS MARCO. Models trained on this dataset show: (1) state-of-the-art performance on advanced reasoning and instruction following datasets; (2) work remarkably well out of distribution due to the ability to respond to user-input prompts; and (3) have explainable reasoning chains that can be given to users or RAG-based systems. Further, we demonstrate that quantized versions of these models retain strong performance while using less compute/memory. Overall, Rank1 shows that test-time compute allows for a fundamentally new type of explainable and performant reranker model for search. [^1] Models, code, and data are available at https://github.com/orionw/rank1
@@ 1 Introduction
Reasoning language models (LMs) like OpenAI’s o1, Deepseek’s R1, and Gemini’s Flash-Thinking have shown improved reasoning abilities through the use of test-time compute, i.e. generating a reasoning chain of tokens that allow the model to “think” before giving the final answer. Another large benefit to these style of models is that the reasoning chain can easily be distilled into smaller models. As shown by Deepseek’s R1 (Guo et al., 2025) smaller models learn incredibly well from simple supervised fine-tuning on the larger model’s reasoning chains.
The benefits that reasoning models bring to general text generation would also be valuable in an information retrieval (IR) context: allowing models additional time to reason why a passage could be relevant, while also allowing an auditable reasoning process to give to the user or RAG system. For this approach to be maximally effective, the model must be able to reason over both query and passage; if applied solely to the query, the reasoning model would not know the passage context and would have to try to infer it. This would be a form of query-expansion (Nogueira et al., 2019) and limits the model’s ability to be precise. Thus, our work focuses on bringing test-time compute to IR in a reranking setting, where the model needs to compute the relevance of an initial top-k candidates.
To accomplish this goal, we sample 635,000 examples of R1’s thought process on the MS MARCO dataset (Nguyen et al., 2016). We then fine-tune a suite of LMs on these reasoning chains and find that they show remarkable reasoning capabilities. Surprisingly, they also exhibit an ability to be prompted despite training from the base LMs only (without instruction fine-tuning) and while having no instruction-based IR training data (only MS MARCO). This includes state-of-the-art performance on the BRIGHT benchmark for reasoning (Su et al., 2024), the NevIR benchmark on complex negation understanding, and the mFollowIR dataset on multilingual instruction-following in IR despite having no non-English reasoning training data (Weller et al., 2025).
Figure 1: Example reasoning traces from R1, used to train Rank1. Note the self-inquisitory reasoning (in blue) where the model questions is if it has the correct answer.
@@ 2.1 Data Preparation
In order to distil from R1, we first need to gather data to use to prompt it. We use the MS MARCO collection due to its diversity in topics and common use in previous work. We use kluster.ai as the API service to access R1 using their batch mode.
We generate data from an equal number (25% of the data) from each of the (1) positive examples in MS MARCO, (2) sampled negatives from Tevatron [^2] From https://huggingface.co/datasets/Tevatron/msmarco-passage-aug (gathered from BM25 and CoCondenser), (3) rank 1-5 hard negatives from mT5-13B, and (4) rank 5-10 hard negatives from mT5-13B. However, we found that R1 classified roughly 80% of the mT5 hard negatives as positives. Thus we did another round of generation using only hard negatives from rank 5-10 and easy negatives from Tevatron. As the mined hard negatives do not have an official label, it is likely that many of them were false negatives and that R1 classified them correctly.
After all generation, our dataset has 635,264 examples of R1 generations, where R1 labeled 62.9% as relevant and 37.1% as non-relevant. We show a plot of these generation lengths in Figure 2 where we see a fairly normal distribution. Although we thought there may be length differences between these four subsets of data, we found that they all had the same rough distribution.
@@ 2.2 Data Mix and Quality Filtering
Since we had a surplus of documented judged relevant, we tried various methods to come to our final data mix. We initially tried using all the data, after balancing for the labels. We found that this performed significantly worse than filtering based on the labels we were most sure about (i.e. the positives from MS MARCO and the negatives from Tevatron). However, even on those subsets, 15% of R1’s final prediction disagreed with the implied labels – thus we filtered out these instances.
Beyond being labeled as positives, we found that a large number of the mT5 mined hard negative samples were noisy. To alleviate this, we used a model trained on the first mix to self-filter the data. [^3] We used the Mistral 24B version as it was the largest that fit on 1 GPU, filtering all instances where the model’s prediction didn’t agree with R1’s prediction. This filtered approximately another 10% of the data, mostly false negatives. Since we still had a surplus of ”relevant” labelled instances, we took all positives from the official MS MARCO positives and all negatives from the self-filtered set. This left us with a training set of 386,336 high quality training samples: 136k from the original MS MARCO positives, 154k from the Tevatron negatives, and 96k from the mT5 negatives. We note that having more negatives is standard: RankLLaMA (Ma et al., 2024) trained on a 15:1 ratio of negatives to positives for 7 million examples.
@@ 2.3 Training
We train three main models from the Qwen 2.5 family of models (Yang et al., 2024). These models have shown improved performance on recent LM benchmarks and have a wide range of models, allowing us to show the effects of scale. We use the 7B, 14B, and 32B parameter models. We show that alternate base models are also effective in Section 4. During initial experiments we found that the base models outperformed their instruction-tuned variants, so we start from the base versions. [^4] This could be because instruct-versions are optimized for chat and math data, whereas we have a large amount of reranking-specific data that doesn’t benefit from chat/math-based instructions.
@@ 3.1 Baselines
We use BM25 (Robertson et al., 1994; Lù, 2024) and mE5-base (Wang et al., 2024) for our first stage models. For reranking models we show mainly other pointwise reranking models (e.g. models that output a score per document): MonoT5-3B (Nogueira et al., 2019), mT5-13B fine-tuned on MMARCO for multilingual tasks (Jeronymo et al., 2023), and RankLLaMA 7 and 13B (Ma et al., 2024). For instruction following tasks we also include FollowIR-7B (Weller et al., 2024a) which was trained solely for instruction-following.
@@
Table 1: The performance of retrieval models on BRIGHT. Numbers for ReasoningRank are taken from their paper (Ji et al., 2024) while BM25 scores are taken from the official BRIGHT paper. All models rerank from the BM25 on GPT-4o CoT top 100 documents, but are not given the GPT-4o CoT. We find a large gap between similar sized rerankers and Rank1 models (sometimes 2x). Bold indicates the best score for that subset in the reranker section.
We show results for listwise models when those scores are available (Pradeep et al., 2023b; Niu et al., 2024; Sun et al., 2023), but we note that they are not comparable – listwise models take an order of magnitude more time at inference due to their sequential dependencies and have the advantage of seeing all documents in their context when reranking. We show them generally as a strong upper bound, as when state-of-the-art (SOTA) LMs (i.e. GPT-4o) are used.
@@ 3.2 Reasoning Capabilities
We show results on the reasoning intensive BRIGHT benchmark (Su et al., 2024) in Table 1. All reranker models judge the top 100 documents found using BM25 on the query plus GPT-4o’s Chain of Thought (CoT) reasoning, which performed significantly better than BM25 without the query expansion (thus including more relevant documents in the top 100). However, at inference time, the rerankers are not given the CoT. We see a large gap between Rank1 and other models: in many cases near double the nDCG@10 score (e.g. 48.8 vs 24.3 for Rank1-7B vs ReasoningRank GPT-4o on Biology, or 18.7 vs 7.7 for Rank1-7B vs RankLLaMA-13B on Stackoverflow).
3.3 Semantic Understanding
@@
Listwise RankGPT 4o-mini 64.1
@@ 3.4 Instruction-Following
We show results on the mFollowIR dataset (Weller et al., 2025) as it illustrates both instruction-following and multilingual capabilities. Table 4 shows results on the cross-lingual setup (En-XX) and Table 4 on the XX-XX task. We have all models rerank the top 100 scores of a strong but small base model mE5-base. We find that Rank1 has much higher nDCG@20 scores (from 0.586 to 0.611 on the multilingual average) compared to the next best model (mT5-13B trained on Multilingual MS MARCO) with 0.517. Furthermore, when considering just instruction following metrics there is a wide gap, especially in the multilingual setting (+10.1 vs +4.8 p-MRR on the custom instruction-trained FollowIR-7B). Other models are closer on the cross-lingual version, but there still remains a notable gap between the best Rank1 and the closest other model (+7.3 vs +3.7 p-MRR).
@@
We also evaluate DL19 scores in Table 5. All models rerank the top 100 passages found using RepLLaMA. We found that our models performed significantly worse on the initial qrels and our analysis quickly showed this was due to the number of unjudged documents that our models ranks higher. As seen in Table 5, the top 10 lists of existing pointwise rerankers is almost entirely judged passages (i.e. 96.1%), whereas Rank1 finds 10-15% less (down to 81.9%).
To remedy this, we annotated all top 10 documents that each model (baseline or Rank1) got wrong or were unjudged. We show examples of incorrectly labeled instances in Appendix C. We found that unjudged documents were mostly relevant, e.g. 70.6% of the unjudged documents for RankLLaMA 14B were relevant and 86.8% were relevant for Rank1-14B, etc. [^5] We release our new judged qrels in the Github above to help facilitate future work. After this annotation fix, we find that Rank1 models are no longer penalized for finding new documents and that performance is generally better than all other models (e.g. 78.6 vs 76.8 nDCG for Rank1-7B vs RankLLaMA 7B).
@@
We show results on BEIR as a comparison in Table 6, using the datasets with less than 2k queries. All models rerank the top 100 documents found using BM25s (Lù, 2024). We find comparable but worse performance with Rank1. Although out of scope for this work, we found a large number of similar issues in BEIR datasets as we do in DL19. We discuss each dataset individually in Appendix F. It appears that traditional reranking benchmarks like DL19 and BEIR – although extremely successful at driving the field forward and still useful for weaker models – are no longer as useful for distinguishing between the best performing rerankers.
@@ 3.6 Test-Time Scaling
Given the success of reasoning language models in scaling test time compute (i.e. getting better results with more tokens used), we also attempted several ways of using extra test-time compute to improve performance. However, using the simple budget-forcing method from s1 (Muennighoff et al., 2025) did not improve performance and actually hurt performance on average. We hypothesize this may be due to the lack of difficulty in the reranking task – after all, reasoning over BRIGHT requires significantly less reasoning than the typical AIME problems used for evaluation. Alternatively, perhaps new techniques are needed to carefully induce the desired results.
@@
Table 6: nDCG@10 results on the BEIR evaluation benchmark. Models rerank the top 100 documents from BM25S (Lù, 2024). Best results on each dataset and the entire benchmark are boldfaced.
@@ 5.1 Advanced Reasoning and Instruction Following in IR
In the last couple years, retrieval systems have started to move beyond simple phrase-based semantic matching, to more complex information retrieval tasks. This has included a focus on new benchmarks: such as reasoning in retrieval (Su et al., 2024), instruction-following capabilities (Weller et al., 2024a; Oh et al., 2024), and retrieval for RAG systems (Lawrie et al., 2024; Mayfield et al., 2024).
On the modeling side, we have seen a surge of interest in models that can understand the meaning behind the user’s query rather than doing phrase-based matching: this includes models like Instructor (Su et al., 2022), TART (Asai et al., 2022), GritLM (Muennighoff et al., 2024), FollowIR (Weller et al., 2024a), Gecko (Lee et al., 2024), and Promptriever (Weller et al., 2024b). These models typically use instruction-based data in their training data, so that they learn to adapt to new user instructions. However, for Rank1 we do not provide any instruction-based training data and fine-tune from the base (non-instruct) version of the LMs – despite this our model shows SOTA ability in these tasks.

### v2

@@ Abstract
We introduce Rank1, the first reranking model trained to take advantage of test-time compute. Rank1 demonstrates the applicability within retrieval of using a reasoning language model (i.e. OpenAI's o1, Deepseek's R1, etc.) for distillation in order to rapidly improve the performance of a smaller model. We gather and open-source a dataset of more than 600,000 examples of R1 reasoning traces from queries and passages in MS MARCO. Models trained on this dataset show: (1) state-of-the-art performance on advanced reasoning and instruction following datasets; (2) work remarkably well out of distribution due to the ability to respond to user-input prompts; and (3) have explainable reasoning chains that can be given to users or retrieval-augmented generation (RAG) systems. Further, we demonstrate that quantized versions of these models retain strong performance while using less compute/memory. Overall, Rank1 shows that test-time compute allows for a fundamentally new type of explainable and performant reranker model for search. [^1] Models, code, and data are available at https://github.com/orionw/rank1
@@ 1 Introduction
Reasoning language models (LMs) like OpenAI's o1, Deepseek's R1, and Gemini's Flash-Thinking have shown improved reasoning abilities through the use of test-time compute, i.e. generating a reasoning chain of tokens that allow the model to ``think" before giving the final answer. Another large benefit to these style of models is that the reasoning chain can easily be distilled into smaller models. As shown by Deepseek's R1 (Guo et al., 2025) smaller models learn incredibly well from simple supervised fine-tuning on the larger model's reasoning chains.
The benefits that reasoning models bring to general text generation would also be valuable in an information retrieval (IR) context: allowing models additional time to reason why a passage could be relevant, while also allowing an auditable reasoning process to give to the user or RAG system. For this approach to be maximally effective, the model must be able to reason over both query and passage; if applied solely to the query, the reasoning model would not know the passage context and would have to try to infer it. This would be a form of query-expansion (Nogueira et al., 2019) and limits the model's ability to be precise. Thus, our work focuses on bringing test-time compute to IR in a reranking setting, where the model needs to compute the relevance of an initial top-k candidates.
To accomplish this goal, we sample 635,000 examples of R1's thought process on the MS MARCO dataset (Nguyen et al., 2016). We then fine-tune a suite of LMs on these reasoning chains and find that they show remarkable reasoning capabilities. Surprisingly, they also exhibit an ability to be prompted despite training from the base LMs only (without instruction fine-tuning) and while having no instruction-based IR training data (only MS MARCO). This includes state-of-the-art performance on the BRIGHT benchmark for reasoning (Su et al., 2024), the NevIR benchmark on complex negation understanding, and the mFollowIR dataset on multilingual instruction-following in IR – despite having no non-English reasoning training data (Weller et al., 2025a).
Figure 1: Example reasoning traces from R1, used to train Rank1. Note the self-inquisitory reasoning (in blue) where the model questions if it has the correct answer.
@@ 2.1 Data Preparation
In order to distil from R1, we first need to gather data to use for prompting it. We use the MS MARCO collection (Nguyen et al., 2016) due to its diversity in topics and common use in previous work. We use kluster.ai as the API service to access R1 using their batch mode with a temperature of 0.3 and a maximum of 1000 output tokens. [^2] In our initial testing we found that all reasoning chains were shorter than 1000 tokens. As each MS MARCO passage is typically around 100-200 token, R1 doesn’t need more than 1000 tokens to reason over it.
We generate data from an equal number (25% of the data) from each of the (1) positive examples in MS MARCO, (2) sampled negatives from Tevatron [^3] From https://huggingface.co/datasets/Tevatron/msmarco-passage-aug (gathered from BM25 and CoCondenser), (3) rank 1-5 hard negatives from mT5-13B, and (4) rank 5-10 hard negatives from mT5-13B. However, we found that R1 classified roughly 80% of the mT5 hard negatives as positives. Thus we did another round of generation using only hard negatives from rank 5-10 and easy negatives from Tevatron. As the mined hard negatives from mT5 do not have an official label, it is likely that many of them were false negatives and that R1 classified them correctly.
After all generation was done, our dataset has 635,264 examples of R1 generations, where R1 labeled 62.9% as relevant and 37.1% as non-relevant. We show a plot of these generation lengths in Figure 2 where we see a fairly normal distribution. Although we thought there may be length differences between these four subsets of data, we found that they all had the same rough distribution.
@@ 2.2 Data Mix and Quality Filtering
Since we had a surplus of documents judged relevant, we tried various methods to arrive at our final data mix. We initially tried using all the data, after balancing for the labels. We found that this performed significantly worse than filtering based on the labels we were most sure about (i.e. the positives from MS MARCO and the negatives from Tevatron). However, even on those subsets, 15% of R1's final prediction disagreed with the implied labels – thus we filtered out these instances.
Beyond being labeled as positives, we found that a large number of the mT5 mined hard negative samples were noisy. [^4] See Appendix Section B where we found that 2/3rds of the mT5 negatives were actually positives. To alleviate this, we used a model trained on the first mix to self-filter the data. [^5] We used the Rank1 Mistral 24B version as it was the largest that fit on 1 GPU, filtering all instances where the model’s prediction didn’t agree with R1’s prediction. This filtered another 10% of the data, mostly false negatives. Since we still had a surplus of "relevant" labeled instances, we took all positives from the official MS MARCO positives [^6] From the MS MARCO small triples file. and all negatives from the self-filtered set. This left us with a training set of 386,336 high quality training samples: 136k from the original MS MARCO positives, 154k from the Tevatron negatives, and 96k from the mT5 negatives. We note that having more negatives than positives is standard: RankLLaMA (Ma et al., 2024) trained on a 15:1 ratio of negatives to positives for 7 million examples.
@@ 2.3 Training
We train three main models from the Qwen 2.5 family of models (Yang et al., 2024). These models have shown improved performance on recent LM benchmarks and have a wide range of models, allowing us to show the effects of scale. We use the 7B, 14B, and 32B parameter models. [^7] We show ablations on smaller Qwen 2.5 models in Section 3.8. We show that alternative base models are also effective in Section 4. During initial experiments we found that the base models outperformed their instruction-tuned variants, so all models are trained from the base version (e.g. no post-training/instruction-training). [^8] This could be because instruct-versions are optimized for chat and math data, whereas we have a large amount of reranking-specific data that doesn’t benefit from chat/math-based instructions.
@@ 3.1 Baselines
We use BM25 (Robertson et al., 1994; Lù, 2024) and mE5-base (Wang et al., 2024) for our first stage models. For reranking models we focus on other pointwise reranking models (e.g. models that output a score per document): MonoT5-3B (Nogueira et al., 2019), mT5-13B fine-tuned on MMARCO for multilingual tasks (Jeronymo et al., 2023), and RankLLaMA 7 and 13B (Ma et al., 2024). For instruction following tasks we also include FollowIR-7B (Weller et al., 2024a) which was trained solely for instruction-following.
@@
Table 1: The performance of retrieval models on BRIGHT. BM25 scores are taken from the official BRIGHT paper. All models rerank from the BM25 on GPT-4o CoT top 100 documents, but are not given the GPT-4o CoT. We find a large gap between similar sized rerankers and Rank1 models (sometimes 2x). Bold indicates the best score for that subset in the reranker section.
We show results for listwise models when those scores are available (Pradeep et al., 2023b; Sun et al., 2023), but note that they are not comparable – listwise models take an order of magnitude more time at inference due to their sequential dependencies and have the advantage of seeing all documents in their context when reranking. We show them generally as a strong upper bound, as when state-of-the-art (SOTA) LMs (i.e. GPT-4o) are used.
@@ 3.2 Reasoning Capabilities
We show results on the reasoning intensive BRIGHT benchmark (Su et al., 2024) in Table 1. All reranker models judge the top 100 documents found using BM25 on the query plus GPT-4o's Chain of Thought (CoT) reasoning, which performed significantly better than BM25 without the query expansion (thus including more relevant documents in the top 100 to test rerankers with). However, at inference time, the rerankers are not given the CoT. We see a large gap between Rank1 and other models: in many cases near double the nDCG@10 score (e.g. 18.7 vs 7.7 for Rank1-7B vs RankLLaMA-13B on Stackoverflow).
@@
Listwise
RankGPT 4o-mini 64.1
@@ 3.4 Instruction-Following
We show results on the mFollowIR dataset (Weller et al., 2025a) as it illustrates both instruction-following and multilingual capabilities. Table 4 shows results on the cross-lingual setup (En-XX) and Table 4 on the XX-XX task. We have all models rerank the top 100 scores of a strong but small base model mE5-base. We find that Rank1 has much higher nDCG@20 scores (from 0.586 to 0.611 on the multilingual average) compared to the next best model (mT5-13B trained on Multilingual MS MARCO) with 0.517. Furthermore, when considering just instruction following metrics there is a wide gap, especially in the multilingual setting (+10.1 vs +4.8 p-MRR on the custom instruction-trained FollowIR-7B). Other models are closer on the cross-lingual version, but there still remains a notable gap between the best Rank1 and the closest other model (+7.3 vs +3.7 p-MRR).
@@
We also evaluate DL19 scores in Table 5. All models rerank the top 100 passages found using RepLLaMA. We found that our models performed significantly worse on the initial relevance judgments (i.e. qrels) and our analysis quickly showed this was due to the number of unjudged documents that our models ranks higher. As seen in Table 5, the top 10 lists of existing pointwise rerankers is almost entirely judged passages (i.e. 96.1%), whereas Rank1 finds 10-15% less (down to 81.9% passages that are judged by humans).
To remedy this, we manually annotated all top 10 documents that each model (baseline or Rank1) got wrong or were unjudged. We show examples of incorrectly or not labeled instances in Appendix D. We found that unjudged documents were mostly relevant, e.g. 70.6% of the unjudged documents for RankLLaMA 14B were relevant and 86.8% were relevant for Rank1-14b, etc. [^9] We release our new judged qrels in the Github above to help facilitate future work. After this annotation fix [^10] We note that ideally one would re-annotate the whole dataset with new pools from many models. This would allow for future models to be able to use the same evaluation. However, given the scale of annotations, this would be out of scope for this work. Thus, we include this experiment mainly to highlight the limitations of DL19 and acknowledge that better evaluations will be needed in future work. Rank1 models are no longer penalized for finding new documents and that performance is generally better than all other models (e.g. 78.6 vs 76.8 nDCG for Rank1-7b vs RankLLaMA 7B).
@@
We show results on BEIR as a comparison in Table 6, using the datasets with less than 2k queries. All models rerank the top 100 documents found using BM25s (Lù, 2024). We find comparable but worse performance with Rank1. Although out of scope for this work, we found a large number of similar issues in BEIR datasets as we do in DL19. We discuss datasets individually in Appendix G. It appears that traditional reranking benchmarks like DL19 and BEIR – although extremely successful at driving the field forward and still useful for weaker models – are no longer as useful for distinguishing between the best performing rerankers.
@@ 3.6 Test-Time Scaling
Given the success of reasoning language models in scaling test time compute (i.e. getting better results with more tokens used), we also attempted several ways of using extra test-time compute to improve performance. However, using the simple budget-forcing method from s1 (Muennighoff et al., 2025) did not improve performance and actually hurt performance in our limited initial experiments. We hypothesize this may be due to the lack of difficulty in the reranking task – after all, reasoning over BRIGHT requires significantly less reasoning than the typical AIME problems used for evaluation. Alternatively, perhaps future work with new techniques are needed to induce the desired results.
@@
Table 6: nDCG@10 results on the BEIR evaluation benchmark. Models rerank the top 100 documents from BM25S (Lù, 2024). Although not in scope for this work, we find that the way many of the datasets in BEIR were constructed limits the ability to judge between the highest performing systems, similar to DL19 (see Appendix G for commentary on individual datasets).
3.7 How much does the reasoning chain help?
To ablate how important the reasoning chain is for Rank1 we train an ablation on the 7B model where the reasoning chain is omitted using the same data and base model. This allows us to isolate the benefits that the reasoning chain itself provides.
We find that our non-reasoning model scores an average of 17.5 on BRIGHT compared to 27.5 with Rank1-7b. This is an improvement over RankLLaMA-7B (13.9) but not the 2x improvement seen by the models with reasoning chains. Thus, we can see that the reasoning chains are what is providing the expressive power that is needed for these complex reranking tasks.
Figure 3: Average nDCG@10 across BRIGHT datasets for various sizes of Rank1. We see that with our training dataset there are diminishing returns after 7B+ parameters and that smaller models still perform very well: Rank1-0.5b scores comparably to RankLLaMA 13B (13.9 average on BRIGHT) despite having 25x less parameters and only slightly more than BERT-sized models.
3.8 Does this work at smaller sizes?
We have previously used 7B+ models, as it is the smallest size typically used for reasoning models applied to other tasks (e.g. AIME, etc.). However, for reranking, is it possible to have strong performance with smaller models?
We follow the same training recipe and train the rest of the Qwen 2.5 family of models (0.5B, 1.5B, 3B) and show the results with the main models in Figure 3. We find that larger models do better (albeit with dimishing returns), but even smaller models do very well comparatively: the 13B RankLLaMA model (13.9 average) only scores slightly higher than the 0.5B version and less than the 1.5B version.
Thus, although test-time compute is typically used with larger models, we see that – at least for reranking – even small models show large benefits.
@@ 5.1 Advanced Reasoning and Instruction Following in IR
In the last couple years, retrieval systems have started to move beyond simple phrase-based semantic matching, to more complex information retrieval tasks. This has included a focus on new benchmarks: such as reasoning in retrieval (Su et al., 2024; Zhao et al., 2024; Xiao et al., 2024), instruction-following capabilities (Weller et al., 2024a; Oh et al., 2024), and retrieval for RAG systems (Lawrie et al., 2024; Mayfield et al., 2024).
On the modeling side, we have seen a surge of interest in models that can understand the meaning behind the user's query rather than doing phrase-based matching: this includes models like Instructor (Su et al., 2022), TART (Asai et al., 2022), GritLM (Muennighoff et al., 2024), FollowIR (Weller et al., 2024a), Gecko (Lee et al., 2024), and Promptriever (Weller et al., 2024b). These models typically use instruction-based data in their training data, so that they learn to adapt to new user instructions. Other works have built multi-agent systems with LLMs that do reasoning (Niu et al., 2024; Ji et al., 2024). However, for Rank1 we use a single model and not provide any instruction-based training data and fine-tune from the base (non-instruct) version of the LMs – despite this our model shows SOTA ability in these tasks.